OpenStudy (anonymous):

IM STUCK HELP!!! Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided. Can someone show me an example?

OpenStudy (anonymous):

@mathsciencehistory

OpenStudy (anonymous):

The remainder is zero.

OpenStudy (anonymous):

@robtobey what?

OpenStudy (anonymous):

binomial is a function of ax^2+bx+c? then i can help, cuz am not used to the englisch names.

OpenStudy (anonymous):

if thats the case its like this, it prob looks like a divition u made when u were about 10 years old but its kinda diffrent: 6x^3+2x^2+24x+8 divided by x^2+4

OpenStudy (anonymous):

write like this: 6 2 24 8 / x^2+0x+4 -6 0 -24 /6x =0 2 0 8/6x+2 0 -2 0 -8 = 0 0 0 0 so our therm in x^3 can be divided by our therm in x^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

was that wat our asking?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

+ i made it abit shroter for typing but the principe is: u write the coefficients of our therm like 6 (x^3) 2(x^2) 24(x) 8 divided (/) x^2+0x+4 now u c what coeff u need to multiply x^2 to get 6x^3 aswer 6x write 6x beneath x^2+0x and use the multiplaction 6x*(x^2+4)= 24x^3 + 24x therefore the -6 and -24 beneath the 6 and 24. after this u need to substract, and thats why i already multiplied it by -1.

OpenStudy (anonymous):

note that its not necesarry u get the same numbers when u do this, but at the end there will be 0 0 0 0 if they can be divided by each otehr.